### v1

# A Roadmap for Software Testing in Open Collaborative Development Environments

Abstract.
Amidst the ever-expanding digital sphere, the evolution of the Internet has not only fostered an atmosphere of information transparency and sharing but has also sparked a revolution in software development practices. The distributed nature of open collaborative development, along with its diverse contributors and rapid iterations, presents new challenges for ensuring software quality. This paper offers a comprehensive review and analysis of recent advancements in software quality assurance within open collaborative development environments. Our examination covers various aspects, including process management, personnel dynamics, and technological advancements, providing valuable insights into effective approaches for maintaining software quality in such collaborative settings. Furthermore, we delve into the challenges and opportunities arising from emerging technologies such as LLMs and the AI model-centric development paradigm. By addressing these topics, our study contributes to a deeper understanding of software quality assurance in open collaborative environments and lays the groundwork for future exploration and innovation.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: International Workshop on Software Engineering in 2030; November 2024; Puerto Galinàs (Brazil)
1. Introduction
Open collaboration is a hallmark of the Internet era in the past ten to twenty years, where geographical distances cease to be a barrier. Platforms like social media networks and online forums have emerged as global hubs for communication, enabling individuals worldwide to exchange ideas, share experiences, and engage in discussions (Gruzd and Haythornthwaite, 2013). Similarly, in the realm of software development, platforms like GitHub have flourished, providing developers from diverse backgrounds with the opportunity to contribute code, discuss technical matters, and address software issues, regardless of their geographic location (Booch and Brown, 2003; Sengupta et al., 2006). As of now, GitHub hosts a vast array of projects spanning various domains, including operating systems, programming languages, and application software, forming a vast and diverse open-source ecosystem (Tsay et al., 2014). Moreover, with the rise of crowdsourced software development, an increasing number of projects leverage large-scale, decentralized, and heterogeneous resources and manpower for development (Lin, 2018; Wang et al., 2022). This trend further expands and deepens the scope of open collaboration.
In such development environments, software quality assurance faces both new challenges and opportunities. The distributed nature of open collaborative development, coupled with diverse contributors and rapid iterations, necessitates novel approaches to ensure the quality and reliability of software products. Additionally, the sheer volume and complexity of open-source projects demand sophisticated quality assurance techniques to effectively manage and mitigate potential risks and issues.
There exist notable roadmaps that have significantly contributed to the understanding and advancement of software quality assurance practices. One such landmark roadmap is “Software Testing: A Research Travelogue (2000–2014),” spearheaded by Professor Alessandro Orso and Gregg Rothermel (Orso and Rothermel, 2014). This seminal work primarily explores software testing techniques and methodologies, offering valuable insights into the evolution of testing technologies over the past decade and a half. By comparison, our focus in this work is on quality assurance, and the factors influencing software quality extend far beyond testing techniques alone. They encompass a myriad of aspects, including process considerations, the dynamics of involved personnel, and more. Particularly, within open collaborative software development paradigms, traditional process techniques may find limited applicability, while human factors become increasingly pronounced. Moreover, while software has long claimed to be a socio-technical system (Mens, 2016), it is within open communities where the true complexity of this socio-technical system’s impact on software quality becomes strikingly apparent. Therefore, a comprehensive overview and roadmap to software quality assurance in open collaborative environments demands a holistic understanding that encompasses not only testing methodologies but also broader contextual factors.
Figure 1. Overview about software quality assurance under open collaborative development environment
Taken in this sense, our exploration into software quality assurance within open collaborative development environments is guided by the recognition that three key dimensions — process, personnel, and technology — form a triad of fundamental factors, influencing software quality. These dimensions represent orthogonal facets that collectively shape the landscape of software development and quality assurance, as visually demonstrated in Figure 1. Each dimension offers unique insights into the challenges and opportunities inherent in ensuring software quality within open collaborative environments. By exploring these three dimensions—process, personnel, and technology—we achieve a holistic grasp of the intricate landscape of software quality assurance within open collaborative development environments. Together, they constitute the foundation of present-day effective methodologies for upholding software quality.
Additionally, we also outline future trends. As emerging technologies like pre-trained Large Language Models (LLMs) continue to evolve, the landscape of software quality assurance research is expected to expand to encompass various aspects. This may include leveraging LLMs for enhancing testing practices, exploring testing methodologies tailored for LLMs and related applications, ensuring quality assurance for auto-generated code produced by AI models, addressing security concerns within AI model ecosystems, and refining process management and collaboration strategies within AI model-centric paradigms. These areas represent promising avenues for future research and innovation in the field of software quality assurance within open collaborative development environments.
The paper is structured as follows: Section 2 to 4 provide an overview of quality assurance researches from the perspectives of process, personnel, and technology, respectively. Section 5 explores the opportunities and challenges in this field, while Section 6 concludes the paper.
2. Process Related
Process, being the cornerstone of software development, encompasses methodologies, workflows, and practices that govern quality assurance in open collaborative environments. It delves into practices like continuous integration, aimed at frequent code integration and automated testing to uphold quality standards. However, challenges lie in optimizing regression testing and prioritizing test cases for timely developer feedback. Additionally, modern code review practices leverage automation and machine learning to streamline the review process and enhance overall code quality. Community involvement further shapes the process dimension, bringing multidimensional knowledge as well as a plethora of noise and redundant information, posing significant challenges to process management.
2.1. Continuous Integration and Modern Code Review
Continuous Integration (CI) is a software development practice designed to encourage development teams to integrate their code frequently into a shared version control repository through automated build and testing processes. Each integration is automatically built and usually validated by regression testing, upon completion of which the developers are provided feedback. The execution of regression tests may require significant computational resources and take hours or even days to be completed. The prolonged execution of regression tests can delay the CI cycles and prevent timely feedback to the developers (Yaraghi et al., 2022). Test case selection and prioritization techniques have been proposed to improve regression testing by selecting and prioritizing test cases in order to provide early feedback to developers. We introduce the Test Case Prioritization (TCP) techniques which share similarities with selection techniques.
Earlier attempts start with heuristics-based techniques, e.g., Elbaum et al. (Elbaum et al., 2014) prioritized tests based on whether they have not been executed for long or have failed in the recent commits, and Haghighatkhah et al. (Haghighatkhah et al., 2018) scheduled tests based on the combination of their previous execution results and similarity. Later, other researchers harness the power of machine learning by using a large amount of historical data in CI, and propose numerous machine learning based TCP techniques which have been demonstrated to be promising. They build neural models to predict the optimal sequence of tests instead of human-defined strategies. In particular, these machine learning based techniques can be categorized into supervised learning-based (Busjaeger and Xie, 2016; Bertolino et al., 2020; Sharif et al., 2021; Yaraghi et al., 2022), and reinforcement learning-based techniques (Spieker et al., 2017; Bagherzadeh et al., 2022; Lima and Vergilio, 2022). The later ones would continuously adjust its prioritization strategy, i.e., it is first tested (i.e., prioritizing the test suite) and then trained based on the prioritization feedback.
Continuous integration and modern code review are complementary practices that work together to ensure code quality, foster collaboration, and drive continuous improvement in software development projects. Modern code review involves systematically reviewing and evaluating proposed changes to the source code of a software project. Reviewers need to thoroughly understand the source code during code review activities and leave review comments to interact with developers, which involves a significant amount of effort and time of reviewers. Researchers are paying attention to automating code review activities, including comment location prediction (Hellendoorn et al., 2021; Shi et al., 2019), review comment recommendation (Gupta and Sundaresan, 2018; Li et al., 2022a, b) and code refinement (Tufano et al., 2021). Specifically, Shi et al. (Shi et al., 2019) proposed the DACE framework based on CNN and LSTM to predict whether a hunk in the code change will be accepted by the reviewers. Hellendoorn et al. (Hellendoorn et al., 2021) used the Transformer architecture to solve this task. To save the time reviewers spent on writing reviews related to common issues such as coding style and documentations, Gupta and Sundaresan (Gupta and Sundaresan, 2018) proposed the LSTM-based model DeepMem. It learns the relations between code changes and review comments, and recommends review comments automatically based on existing code reviews and code changes. CodeReviewer (Li et al., 2022a) and AUGER (Li et al., 2022b) utilize the pre-training models for the automatic generation of review comments. Tufano et al. (Tufano et al., 2021) used deep learning techniques to automate a different task in code review. They train a Transformer model to revise the contributors’ code to implement the requirements from the review comments.
2.2. Community Involvement and Information Filtering
Diverse individuals with different roles and identities come together to contribute towards open source software development. This collaborative effort encompasses developers, testers, designers, project managers, and end-users, among others, who collectively engage in activities such as issue reporting, code review, and discussions to drive the improvement and refinement of software products. However, as participation in these collaborative platforms grows, it inevitably leads to the generation of a vast amount of information, including issue reports, discussions, and contributions. While this influx of information can provide valuable insights and perspectives, it also poses challenges related to information overload, redundancy, and quality assurance. To address these challenges, various information filtering techniques have emerged within open collaborative development environments. These techniques aim to streamline the information flow, improve signal-to-noise ratio, and enhance the overall quality of contributions.
Duplicate detection is the most-commonly employed technique, which aims at identifying and eliminating redundant or duplicate information, such as duplicate issue reports or discussions on similar topics. For example, Nguyen et. al. (Nguyen et al., 2012) applied information retrieval techniques for duplicate detection by computing the textual similarity between two reports. Sun et. al. (Sun et al., 2010) designed a set of features for measuring the reports’ similarity in terms of textual descriptions and attributes, and employed machine learning techniques for duplicate detection. Yang et. al. (Yang et al., 2016) modeled the semantic similarity of reports using deep learning techniques for duplicate detection. In addition, Huang et. al. (Huang et al., 2020) and Zhang et. al. (Zhang et al., 2023a) respectively conducted experimental evaluations of the commonly-used approaches for duplicate detection.
Knowledge extraction from developer conversations also plays an increasingly significant role for accelerating the development process. Di Sorbo et al. (Sorbo et al., 2015) proposed a taxonomy of intentions to classify sentences in developer mailing lists. Qu et al. (Qu et al., 2019) utilized machine learning methods to perform user intent prediction. Chowdhury et al. (Chowdhury and Hindle, 2015) filtered out off-topic discussions in programming IRC channels by engaging Stack Overflow discussions as positive examples and YouTube video comments. Shi et al. (Shi et al., 2021) extracted the issue-solution pairs from massive chat messages to facilitate OSS developers to check and fix issues.
3. Personnel Related
In the realm of open-source software development, the involvement and collaboration of individuals play a crucial role in shaping the quality and success of projects. This dimension delves into various aspects concerning individuals, including code review practices, worker recommendation for issues and testing tasks, and support for newcomers.
3.1. Code Reviewer Recommendation
Modern code review involves two roles - code contributors and code reviewers. In a simplified code review procedure, code contributors commit code changes and propose a code review request; code reviewers respond to the request asynchronously, then review code changes and give feedback. However, practitioners often find reviewers manually or, at best, semi-manually in practice. Therefore, researchers proposed code reviewer recommendation approaches to recommend code reviewers efficiently. These approaches are achieved by learning the historical data of code review activity, including the developers’ code changes and participation in prior code reviewers (Chen et al., 2022; Hu et al., 2020; Rong et al., 2022). For example, Thongtanunam et al. (Thongtanunam et al., 2015) leverages file paths for expertise-based reviewer recommendation, which assumes that the files located in close files may share similar functionality and are likely to be reviewed by reviewers with common experience. Zanjani et al. (Zanjani et al., 2015) considers the reviewing history (review number, review time). It counts the number of comments to the file as part of scores. It also considers the frequency and recency of historical contributors to measure the reviewer expertise and give a final recommendation. Motivated by the fact that a revision may have multiple reviewers and they may impact each other through publicly posted comments, Rong et al. (Rong et al., 2022) adopted the hypergraph technique to model this high-order relationship and developed HGRec.
3.2. Issue Triaging
Various issues appear during software development and maintenance, and issue fixing is a time-consuming and costly task. Once a issue report is received, assigning it to a suitable developer within a short time interval can reduce the time and cost of the issue fixing process. This assignment process is known as issue triaging. Issue triaging is a time-consuming process since often a large number of developers are involved in software development and maintenance. To aid in finding appropriate developers, earlier practice adopted techniques as machine learning, graph analysis, fuzzy set, and topic modeling. For example, Anvik et al. (Anvik et al., 2006) utilized machine learning methods to solve it. Jeong et al. (Jeong et al., 2009) proposed to use a bug tossing graph to improve issue triaging prediction accuracy. Tamrawi et al. (Tamrawi et al., 2011) proposed a method called Bugzie, which uses a fuzzy set and cache-based approach to increase the accuracy of issue triaging. Xia et al. (Xia et al., 2017) proposed a specialized topic modeling algorithm named multi-feature topic model for issue triaging. Later studies used deep learning for bug triaging, e.g., Lee et al. (Lee et al., 2017) applied word embedding to train a CNN based classifier. Dipongkor et al. (Dipongkor and Moran, 2023) conducted the experimental evaluation about fine-tuning the transformer-based language models for this task.
3.3. Crowd Tester Recommendation
With the emergence of crowdsourced software development, more and more projects are utilizing extensive, decentralized, and diverse resources and manpower for their development efforts. Among them, a significant category is crowdsourced software testing (also short for crowdtesting). Unlike typical software development projects, testing tasks typically require a group of testers, ideally with diverse backgrounds, to achieve the goal of diversified testing and covering different areas of the software. Therefore, different from the worker recommendation for general crowdsourcing tasks, the tester recommendation for crowdtesting tasks tend to take into account the diversity of the recommended workers. For example, Wang et al. (Wang et al., 2021) propose a multi-objective crowd tester recommendation approach, which aims at recommending crowd tester by maximizing the bug detection probability of testers, the relevance with the test task, the diversity of testers, and minimizing the test cost. Additionally, previous studies on worker recommendation mainly focus on one-time recommendations with respect to the initial context at the beginning of a new task. However, for crowdtesting, a typical task can last from 3 days to 2 weeks, during when crowd testers can freely conduct the testing and submit reports. Taking this into account, Wang et al. (Wang et al., 2020) point out the need for accelerating crowdtesting by recommending appropriate testers in a dynamic manner, and they propose a context-aware in-process crowd testers recommendation approach, to detect more bugs earlier and potentially shorten the testing period.
3.4. Support for New Comers
A continuous influx of newcomers is essential for the survival, long-term success, and continuity of OSS projects [1, 8]. However, researches pointed out that the lack of appropriate tasks, up-to-date documents, new tools are the factors that might hinder developers from joining the project (Shibuya and Tamai, 2009; Zhou and Mockus, 2012). To help newcomers find their tasks, Huang et al. (Huang et al., 2021b) defined 79 features to characterize the suitable tasks, i.e., good first issue (GFIs), and built machine learning models to predict the GFIs with the proposed features. Similarly, Xiao et al. (Xiao et al., 2022) modeled an issue with features from multiple dimensions (content, background, and dynamics) and built the classifier to generate its probability of being a GFI. Following that, Tan et al. (Tan et al., 2023) investigated whether it is enough to recommend tasks to newcomers. They analyze the resolution process of GFIs, and find that expert involvement positively correlates with newcomers’ successful contributions. Canfora et al. (Canfora et al., 2012) focused on identifying and recommending mentors in software projects by mining data from mailing lists and versioning systems.
4. Technology Related
In the realm of technology, advancements in software quality assurance are driven by innovative approaches and tools tailored to the demands of open collaborative development environments. From ensuring the integrity of third-party libraries to leveraging crowd intelligence for testing, technological solutions are essential for maintaining software quality amidst the complexities of collaborative development. Through automated identification of vulnerable libraries and harnessing crowd-contributed data for testing enhancements, these technological interventions aim to fortify software ecosystems and enhance overall quality assurance practices. Another significant change has been the rise and prevalence of Artificial Intelligence (AI) algorithms and applications, which has spurred the development of various testing techniques for AI models.
4.1. Quality Assurance for Software Ecosystem
Third-party libraries play a key role in software development because they allow developers to reuse common functionalities instead of reinventing the wheel and substantially improve the productivity of developers. However, outdated third-party libraries can be commonly used but seldom updated, or updated in a delayed way, which might cause software bugs or increase attack surfaces in client projects. Huang et al. (Xi et al., 2019; Huang et al., 2022) conducted large-scale analyses to characterize usages, updates and risks of third-party libraries in Java open-source projects, and proposed a security bug-driven altering system to assist developers for making confident decisions about third-party library version updates. There are studies focusing on automating the identification and updating of deprecated third-party libraries, i.e., APIs. Some of them employed pattern-based techniques, i.e., they rely on the replacement messages in API documents to derive the update patterns, which are subsequently used to guide the replacement of the deprecated APIs (Huang et al., 2021a). Others proposed automating API updates by mining update patterns from the code change history (Fazzini et al., 2019; Haryono et al., 2021). Nevertheless, this source of information only covers a small portion of the deprecated APIs, and is usually delayed.
Software composition analysis is also widely adopted to identify vulnerable libraries used in software applications. For example, Dong et al. (Dong et al., 2019) leveraged named entity recognition to recognize affected libraries from vulnerability description. Chen et al. (Chen et al., 2020) and Lyu et al. (Lyu et al., 2023) formulated this affected library identification problem as an extreme multi-label learning problem, which use library coordinates as labels, and predict labels for each vulnerability based on features extracted from vulnerability description and direct references. Wu et al. (Wu et al., 2024) identified affected libraries and their ecosystems for vulnerabilities via a learning-to-rank technique.
4.2. Testing Benefited from Crowd Intelligence
In the context of open collaborative software development, a vast amount of data contributed by developers with diverse backgrounds is aggregated. This data encapsulates rich knowledge about software quality assurance, and can be harnessed to empower the testing techniques. For example, Mao et al. (Mao et al., 2017) extracted the test scripts from crowdbased testing to automatically infer the reusable high-level event sequences for enhancing the automated mobile testing. Liu et al. (Liu et al., 2021) collected a large number of GUI screenshots with UI display issues from the crowdsourced testing platform, and used them to train a visual understanding model for automatically detecting the GUIs with display issues and locating the buggy region. Wei et al. (Wei et al., 2022) mined the code/models from open source to obtain the code snippets from the deep learning library documentation, library developer tests, and deep learning models in the wild, then leveraged these information to perform fuzz testing for deep learning libraries.
Figure 2. Distribution of testing tasks with LLMs (Wang et al., 2024b)
In recent years, the pervasive advancements in pre-trained Large Language Models (LLMs) have profoundly impacted various domains, including software testing. These models have been increasingly harnessed to bolster testing capabilities across different facets, ranging from enhancing test coverage in unit testing to diversifying test case generation in integration testing. There is a relevant literature review titled “Software Testing with Large Language Models: Survey, Landscape, and Vision” (Wang et al., 2024b), which provides a comprehensive overview of the utilization of LLMs in software testing. It analyzes 102 relevant studies that have employed LLMs for software testing, examining them from both the software testing and LLMs perspectives. As demonstrated in Figure 2, LLMs are commonly used for test case preparation (including unit test case generation, test oracle generation, and system test input generation), program debugging, and bug repair, while there is currently no practices for applying LLMs in the tasks of early testing life-cycle (such as test requirement, test plan, etc).
For unit test case generation, a majority of the earlier published studies adopt the pre-training or fine-tuning schema, for example Alagarsamy et al. (Alagarsamy et al., 2023) first pre-trained the LLM with the focal method and asserted statements to enable the LLM to have a stronger foundation knowledge of assertions, then fine-tuned the LLM for the test case generation task where the objective is to learn the relationship between the focal method and the corresponding test case. By comparison, most later studies typically focus on how to design the prompt, to make the LLM better understanding the context of the task. Yuan et al. (Yuan et al., 2023) performed an empirical study to evaluate ChatGPT’s capability of unit test generation with both a quantitative analysis and a user study in terms of correctness, sufficiency, readability, and usability. And results show that the generated tests still suffer from correctness issues, including diverse compilation errors and execution failures. They then proposed an approach that leveraged the ChatGPT itself to improve the quality of its generated tests with an initial test generator and an iterative test refiner. Specifically, the iterative test refiner iteratively fixed the compilation errors in the tests generated by the initial test generator, which follows a validate-and-fix paradigm to prompt the LLM based on the compilation error messages and additional code context.
For system test input generation, it varies for specific types of software being tested. For example, for mobile applications, the test input generation requires providing a diverse range of text inputs or operation combinations (e.g., click a button, long press a list) (Liu et al., 2023a, 2024), which is the key to testing the application’s functionality and user interface; while for Deep Learning (DL) libraries, the test input is a program which covers diversified DL APIs (Deng et al., 2023b, a). For example, Liu et al. (Liu et al., 2024) formulates the test input generation of mobile GUI testing problem as a Q&A task, which asks LLM to chat with the mobile apps by passing the GUI page information to LLM to elicit testing scripts (i.e., GUI operation), and executing them to keep passing the app feedback to LLM, iterating the whole process. The proposed GPTDroid also introduces a functionality-aware memory prompting mechanism that equips the LLM with the ability to retain testing knowledge of the whole process and conduct long-term functionality-based reasoning to guide exploration. Deng et al. (Deng et al., 2023b) used both generative and infilling LLMs to generate and mutate valid/diverse input DL programs for fuzzing DL libraries. In detail, it first uses a generative LLM (CodeX) to generate a set of seed programs (i.e., code snippets that use the target DL APIs). Then it replaces part of the seed program with masked tokens using different mutation operators and leverages the ability of infilling LLM (InCoder) to perform code infilling to generate new code that replaces the masked tokens.
4.3. Testing for AI Models and Applications
The rise of Artificial Intelligence (AI) applications raises concerns about trustworthiness, particularly in safety-critical domains such as self-driving systems and medical treatments. Software testing plays a crucial role in detecting and addressing discrepancies between expected and actual behaviors in these applications. However, testing AI systems presents unique challenges due to their statistical nature, evolving behavior, and the oracle problem.
Figure 3. Details of machine learning testing (Zhang et al., 2022a)
The notable systematic review titled “Machine Learning Testing: Survey, Landscapes and Horizons” (Zhang et al., 2022a) presents an extensive examination of methodologies for assessing machine learning (including deep learning) systems. It encompasses 144 papers that explore various aspects of testing properties (such as correctness, robustness, and fairness), testing components (including data, learning programs, and frameworks), workflows (encompassing test generation and evaluation), application scenarios (such as autonomous driving and machine translation), as shown in Figure 3.
Adversarial inputs represent a critical concept within AI testing, as they play a significant role in assessing the robustness of AI models. These inputs are deliberately perturbed based on the original inputs, often deviating from the typical data distribution encountered in real-world scenarios. By subjecting models to these carefully crafted inputs, testers can identify potential weaknesses and shortcomings of AI models. For example, Zhou et al. (Zhou et al., 2020) proposed DeepBillboard to generate realworld adversarial billboards that can trigger potential steering errors of autonomous driving systems. Sun et al. (Sun et al., 2020) automatically generate test inputs via mutating the words in translation inputs for testing machine translation systems. In order to generate translation pairs that ought to yield consistent translations, their approach conducts word replacement based on word embedding similarities.
The test oracle problem is challenging, because many machine learning algorithms are probabilistic programs. Metamorphic relations are widely studied to tackle the oracle problem, and they are based on transformations of training or test data that are expected to yield unchanged or certain expected changes in the predictive output. For example, Dwarakanath et al. (Dwarakanath et al., 2018) applied metamorphic relations to image classifications with SVM and deep learning systems. The changes on the data include changing the feature or instance orders, linear scaling of the test features, normalisation or scaling up the test data, or changing the convolution operation order of the data. Tian et al. (Tian et al., 2018) and Zhang et al. (Zhang et al., 2018) stated that the autonomous vehicle steering angle should not change significantly or stay the same for the transformed images under different weather conditions.
Test adequacy evaluation aims to discover whether the existing tests have a good fault-revealing ability. It provides an objective confidence measurement on testing activities. In traditional software testing, code coverage measures the degree to which the source code of a program is executed by a test suite (Zhang et al., 2019). Unlike traditional software, code coverage is seldom a demanding criterion for AI testing, since the decision logic of an AI model is not written manually but rather it is learned from training data. Pei et al. (Pei et al., 2017) proposed the first coverage criterion, neuron coverage, particularly designed for deep learning testing. Neuron coverage is calculated as the ratio of the number of unique neurons activated by all test inputs and the total number of neurons in a DNN. In particular, a neuron is activated if its output value is larger than a user-specified threshold. Following that, there are more fine-grained criteria, like k-multisection neuron coverage, neuron boundary coverage, and strong neuron activation coverage, etc.
With the advancement of LLMs, there are also studies focusing on the robustness evaluation of LLMs. For example, Wan et. al. (Wan et al., 2024) developed an approach for triggering logical reasoning failures in LLMs. Wang et. al. (Wang et al., 2024a) proposed to automatically detect the issues related with social bias in image generation models.
5. Challenges and Opportunities
Despite the advancements and breakthroughs discussed in the preceding sections, open collaborative development environments continue to present numerous challenges, which suggests ongoing research opportunities and emerging trends in the coming years.
5.1. Leveraging LLMs for Enhancing Testing
Although software testing with LLMs has undergone significant growth in the past three years, it is still in its early stages of development, and numerous challenges and open questions need to be addressed.
The first is the challenge for achieving high coverage. Exploring the diverse behaviors of the software under test to achieve high coverage is always a significant concern in software testing. In this context, test generation differs from code generation, as code generation primarily focuses on producing a single, correct code snippet, whereas software testing requires generating diverse test inputs to ensure better coverage of the software. Although setting a high temperature can facilitate the LLMs in generating different outputs, it remains challenging for LLMs to directly achieve the required diversity. For example, for unit test case generation, in SF110 dataset, the line coverage is merely 2% (Siddiq et al., 2023). For system test input generation, in terms of fuzzing DL libraries, the API coverage for TensorFlow is reported to be 66% (2215/3316) (Deng et al., 2023b). Current studies typically utilize mutation testing together with the LLMs to generate more diversified outputs. For example, when fuzzing a DL library, instead of directly generating the code snippet with LLM, Deng et al. (Deng et al., 2023b) replace parts of the selected seed (code generated by LLM) with masked tokens using different mutation operators to produce masked inputs. They then leverage the LLM to perform code infilling to generate new code that replaces the masked tokens, which can significantly increase the diversity of the generated tests. Other potential research direction could involve utilizing testing-specific data to train or fine-tune a specialized LLM that is specifically designed to understand the nature of testing. By doing so, the LLM can inherently acknowledge the requirements of testing and autonomously generate diverse outputs.
The second is the challenge in test oracle problem. The oracle problem has been a longstanding challenge in various testing applications, e.g., testing machine learning systems (Zhang et al., 2022b) and testing deep learning libraries (Deng et al., 2023b). To alleviate the oracle problem to the overall testing activities, a common practice is to transform it into a more easily derived form, often by utilizing differential testing (Sun et al., 2023) or focusing on only identifying crash bugs (Liu et al., 2024). Exploring the use of LLMs to derive other types of test oracles represents an interesting and valuable research direction. Specifically, metamorphic testing is also widely used in software testing practices to help mitigate the oracle problem, yet in most cases, defining metamorphic relations relies on human ingenuity. Another promising avenue is exploring the capability of LLMs to automatically generate test cases based on metamorphic relations, covering a wide range of inputs. The advancement of multi-model LLMs like GPT-4 may open up possibilities for exploring their ability to detect bugs in software user interfaces and assist in deriving test oracles. By leveraging the image understanding and reasoning capabilities of these models, one can investigate their potential to automatically identify inconsistencies, errors, or usability issues in user interfaces.
The third is the challenge in the real-world application of LLMs in software testing. Due to concerns regarding data privacy, when considering real-world practice, most software organizations tend to avoid using commercial LLMs and would prefer to adopt open-source ones with training or fine-tuning using organization-specific data. Furthermore, some companies also consider the current limitations in terms of computational power or pay close attention to energy consumption, they tend to fine-tune medium-sized models. It might be quite challenging for these models to achieve similar performance to what existing papers have reported. Recent research has highlighted the importance of high-quality training data in improving the performance of models for code-related tasks [152], yet manually building high-quality organization-specific datasets for training or fine-tuning is time-consuming and labor-intensive. To address this, one is encouraged to utilize the automated techniques of mining software repositories to build the datasets, for example, techniques like key information extraction techniques from Stack Overflow (Shi et al., 2019) offer potential solutions for automatically gathering relevant data. In addition, exploring the methodology for better fine-tuning the LLMs with software-specific data is worth considering because software-specific data differs from natural language data as it contains more structural information, such as data flow and control flow. Previous research on code representations has shown the benefits of incorporating data flow (Guo et al., 2021).
5.2. Testing for LLMs and Related Applications
There have been numerous research efforts on testing machine learning and deep learning models, however, in the context of general artificial intelligence, i.e., LLMs, there has been relatively less exploration. Example researches include robustness evaluation of LLMs in logical reasoning tasks (Wan et al., 2024), evaluations of LLMs in task automation (Shen et al., 2023), heterogeneous value alignment evaluation of LLMs (Zhang et al., 2023b), etc. However, there are still many other dimensions of capabilities in LLMs that need to be tested and evaluated, such as contextual understanding and situational comprehension. Additionally, there is a lack of systematic testing methodology specifically designed for LLMs. Furthermore, with the emergence of multi-modal LLMs, there is a need for more research and attention on testing such models.
Furthermore, LLMs are not only utilized in isolation but are also integrated into applications, with a typical example being the LLM-as-agent paradigm (Xi et al., 2023; Liu et al., 2023b). In this scenario, LLMs engage in cognitive understanding, decision-making, and other tasks to support specific application scenarios. This integration extends beyond just the LLMs itself, and encompasses various external factors. For instance, it involves external environments, memory modules for storing historical information from interactions, external knowledge bases for retrieving the latest information, and tools or external services for utilization. Ensuring the robustness and reliability of LLM-as-agent systems requires testing not only the individual components but also their integration and interaction. This involves verifying the model’s ability to effectively leverage external resources, adapt to dynamic environments, and maintain coherence across diverse tasks and contexts. Additionally, testing strategies should encompass scenarios that stress the system’s capabilities, such as handling unexpected inputs, responding to novel situations, and gracefully recovering from errors or disruptions. By adopting a holistic testing approach, developers can gain confidence in the functionality and performance of LLM-as-agent systems across various real-world scenarios.
GPTs, as a new form of service based on LLMs, will make the GPT Store a new channel for people to access applications, similar to the Google Play Store or the Apple App Store in the era of mobile applications. They are custom versions of GPT tailored for specific purposes, allowing users to create personalized iterations of GPT to better suit their needs (OpenAI, 2024). These customized GPTs can be designed for various tasks, such as teaching children math, providing assistance in board games, or generating stickers. Therefore, ensuring the quality and reliability of all GPTs on the platform is a critical concern (Zhang et al., 2024). For example, there may be instances where certain GPTs engage in illegal activities such as unauthorized access to user privacy. Effective testing mechanisms must be established to address such concerns. Moreover, discrepancies between the descriptions of GPT capabilities and their actual performance may arise, necessitating robust testing procedures to ensure consistency and accuracy. Additionally, leveraging user-contributed feedback and reviews on the GPT Store platform could serve as a valuable resource for identifying defects and improving GPT performance.
5.3. Quality Assurance for Auto-generated Code
With the advancement of LLMs and their remarkable performance in code generation tasks, developers are increasingly relying on these models for various coding and debugging tasks. Recently, there have been reports of groundbreaking developments, such as Microsoft’s creation of an AI programmer named AutoDev (Microsoft, 2024), capable of mastering full-stack skills, which can not only write code and debug, but also train models and even bid for projects on the largest freelancing platform, Upwork. As a result, the proliferation of AI-generated code inevitably finds its way into the open-source community, presenting significant challenges for ensuring the overall quality of the open-source ecosystem.
The integration of AI-generated code into open-source projects raises concerns regarding code reliability, maintainability, and security. While AI models demonstrate impressive capabilities in code generation, there remain uncertainties about the robustness and correctness of the generated code (Ouyang et al., 2023). Additionally, the lack of human oversight in the code generation process may lead to the introduction of vulnerabilities or inefficiencies that could compromise the integrity of software systems. Furthermore, the rapid adoption of AI-generated code in open-source projects exacerbates the challenge of maintaining code quality standards across diverse programming languages and development environments. The sheer volume of AI-generated contributions necessitates scalable quality assurance processes to ensure that the code meets established coding standards, adheres to best practices, and complies with project-specific requirements.
To address these challenges, it becomes imperative to develop comprehensive quality assurance frameworks tailored specifically for AI-generated code. Such frameworks should encompass automated code review mechanisms, static analysis tools, and robust testing methodologies to detect and rectify potential issues in the generated code. Moreover, fostering collaboration between AI developers, software engineers, and open-source maintainers is essential to facilitate the seamless integration of AI-generated contributions while upholding the quality and reliability of open-source software projects.
5.4. AI Models Ecosystem and Security
In addition to platforms like GitHub, which host a plethora of code contributions from programmers worldwide, the advancement of AI technology has elevated AI models to crucial components in software development. Many developers now rely on pre-trained AI models downloaded from repositories or fine-tune these models with domain-specific data to achieve specific outcomes. This trend has spurred the emergence of platforms such as Hugging Face (Face, 2024), which serves as repositories for AI models and facilitate their distribution and usage.
In such platforms, where AI models are distributed and utilized, various general security risks may arise. For example, data poisoning can inject manipulated data during model fine-tuning, potentially leading to biased outcomes (Shafahi et al., 2018), while backdoor attacks entail covertly implanting vulnerabilities into AI models, compromising their integrity (Liu et al., 2018). Ensuring the quality of public AI models hosted on these platforms is essential to foster trust and reliability among users and promote the widespread adoption of AI technology. This includes secure data handling practices to protect privacy, implementing mechanisms for model authentication and integrity verification, and deploying defense mechanisms against adversarial attacks, such as robust model testing and validation techniques. Moreover, fostering a culture of security awareness and promoting collaboration between security experts and AI practitioners can help identify and address potential vulnerabilities effectively.
5.5. Process Management and People Collaboration in AI Model-centric Paradigm
The emerging trend of utilizing AI models, downloaded, fine-tuned, and shared on platforms like Hugging Face, may necessitate additional forms of process management and collaboration among individuals to effectively support the development workflow.
In such a scenario, traditional process management methodologies may face challenges in adapting to the dynamic and decentralized nature of AI model-centric open collaboration. One potential challenge is maintaining version control and ensuring consistency across different iterations of AI models, especially when multiple contributors are involved in fine-tuning and uploading. Additionally, ensuring the quality and reliability of AI models may require new strategies for testing, validation, and integration into existing systems. Moreover, effective collaboration within the context of AI model-centric open collaboration may also face challenges. One potential challenge arises from coordinating efforts across diverse contributors with varying levels of expertise and objectives specific to AI model fine-tuning, validation, and integration. This complexity can lead to issues such as conflicting modifications, redundant work, or difficulties in tracking contributions, which are particularly pertinent in the context of AI model development where nuances in algorithmic implementations and data processing can significantly impact performance and outcomes.
Potential strategies for process management and collaboration could include implementing robust version control systems tailored to AI model development, establishing clear guidelines and best practices for contributors. Additionally, harnessing collaborative tools and platforms tailored specifically to AI model-centric development, equipped with functionalities optimized for tasks such as algorithmic validation, model versioning, and data management, could significantly enhance coordination and communication among contributors.
6. Conclusion
In the modern digital landscape, open collaborative software development environments have emerged as a significant trend in the field of software engineering. This paper has explored various dimensions of quality assurance within these environments, encompassing process, personnel, and technology aspects. It also analyze the challenges and opportunities with the emerging of LLMs, and looking ahead, there would be more testing techniques empowered by LLMs, as well as testing techniques tailored for LLMs and related scenarios.